Project

General

Profile

10-Sep-2020

Rick -- Been working with Simon on CRV testing, some progress and issues. Has not had a chance to test 4 repaired DTCs.

Simon -- Reading out noise, DTC is sending out 10K heartbeats, and data requests, then data goes through DTC and Rick provides dump of data packets. Byte count mismatch currently, believe it is an issue with packet generator in CRV ROC. Can get about 24 hits per event on average within latency requirement (<100us). Delay seems to be 30us + 3us * hits. Will try more FEBs soon. Sten thought the fixed delay was 80us. We can maintain 24 hits on average per FEB.

Lei -- Gary is asking if AD torture file causes loss at FEBs? This is on todo list. Still need leverage PREFETCH packet. ON spill should be more than 24 hits.

100:1 trigger rate ⇒ 170 us period
200us * .4/1.4 + 100us *1/1.4 = 129 us Steady State.

Torture sim did not have correct CRV geometry. Might be a factor of 2 off. CRV CRR hopefully before next IPR.

Antonio - completed XY chart widget (2 PVs) and picture with mapped links widget (tool is imagemap.net). Starting with Franco to explore Calo ROC interface to otsdaq. Will work with DIRAC in Pisa.

Vivian -- Working on Tracker test stand. Moved daq09 to lab3. Getting EPICS variables over to daq15. Will attempt to archive the data. Will attempt to connect a panel to DTC next week. Evan reproduced issues with DQM that Vivian had and will work to fix it. Still do not have PVs from DTC,.. Antonio needs to look at file access solution.

Bertrand -- Still thinking about optimization
Ron -- Eric is working on new release. There was an art bug, so Eric is making another release the next days. Thinking about TRACE delimiter /src /src*, could go to next slash.

Greg -- how do we monitor deadtime? Perhaps we should merge a minimum set of data from each source at the DQM level to accomplish this. October 21-24 is Collaboration meeting.

Gary --
Karen --
Ryan -- Iris is down to 1 day per week. We have 10 DTCs in a daisychain, 9 were able to get locked on Friday, by Monday only 7 had stayed locked. The fanout module is functional, need to hook it up.